# Information theory for data-driven model reduction in physics and biology

### v3

@@ III Information theoretic quantities for dynamical systems
Figure S2: Information gain due to an eigenvalue pile-up. (a) To study the role of a gap closing compared to the pile-up of eigenvalues (beyond the dominant one) in the mutual information $I(X_{t};X_{t+\Delta t})$ we build a matrix $P$ with which we can tune the eigenvalues via a parameter $\varepsilon$. (b) Eigenvalues for the matrix shown in (a) with varying parameter values $\varepsilon$; for $\varepsilon\rightarrow 0.5$, the spectral gap $\lambda_{0}-\lambda_{1}$ closes. (c) Eigenvalues for the matrix shown in (a), but with all $\frac{1}{2}(\frac{1}{2}-\varepsilon)^{2}$ terms subtracted with a constant (here $1/8$). Here, the spectral gap $\lambda_{0}-\lambda_{1}$ does not close but there is an accumulation of eigenvalues at $\varepsilon\rightarrow 0.5$. (d-e) Mutual information for $\Delta t=1$ for the matrices in (b-c), respectively, as a function of $\varepsilon$. The curve from the other scenario is shown in gray for comparison. Both exhibit a clear peak in information which differs only slightly in magnitude, showing that the information peak in Fig. 2d-f of the main text is not due to the closing gap alone, but rather due to contributions from subdominant eigenvalues.
@@
$D_{\text{KL}}\left(p(x_{n+1}|x_{n},x_{n-1})\lVert p(x_{n+1}|x_{n})\right).$ (S35)
@@
A system which is non-Markovian can be made Markovian by introducing time-lagged variables, which follows from Taken’s embedding theorem [104]. One constructs a new dynamical variable which is a concatenation of the previous states of the system,
@@ IV Derivations
In Ref. [38] (and reproduced above in (S5)), it is shown that the encoding minimizing the IB objective satisfies the implicit equation
@@
In the $D_{\text{KL}}$ in (S48), the integration is done with respect to the common random variable $X_{t+\Delta t}$ of the two conditional distributions.
@@
Note the appearance of the encoder $p(h_{t}|x_{t})$, making (S48) only an implicit solution of the optimization problem.
@@
The Kullback-Leibler divergence in (S48) can be represented as a sum of two terms,
@@
The first term has no dependence on $h_{t}$ and can hence be absorbed into the normalization $N(x)$. Plugging the decomposition (S52) into the second term leads directly to
@@
IV.1 Perturbation theory at the uniform encoder
@@
Mathematical result – Consider a transfer operator $U$ with infinitesimal generator $\mathcal{L}_{U}$. Assume that $\mathcal{L}_{U}$ has a discrete spectrum with eigenvalues satisfying $0=\text{Re }\lambda_{0}>\text{Re }\lambda_{1}>\text{Re }\lambda_{2}\gg\text{Re }\lambda_{3}$.... Then, for $\beta$ just above the first IB transition $\beta_{1}$ so that $f_{n}(h)\rightarrow 0$, the optimal encoder is given approximately by
@@
To show this we compute the $f_{n}(h)$ coefficients at the onset of instability. The instability at the first IB transition at $\beta=\beta_{1}$ corresponds to the emergence of negative eigenvalues in the Hessian of the IB loss,
@@
Details of the calculations leading to these results are provided in the appendix. Angled brackets with no subscript correspond to an average with respect to the steady state distribution, $\langle\cdot\rangle=\int\text{d}x\,\cdot\,p(x)$. Written out more explicitly, the second term is
@@
which are increasing in order (remember $\lambda_{i}\leq 0$). It follows that at the first transition, only $f_{1}$ becomes non-zero, and hence the encoder takes the form given by (S47). In the equilibrium case, the encoder learns exclusively the first eigenfunction, with no correction due to the second eigenfunction.
@@
Similar to the equilibrium case, we see that at the first transition the encoder depends only on $f_{1}$, giving (S47). In contrast to the equilibrium case, there may be a small correction due to the second eigenfunction, however this becomes exponentially small for long times $\Delta t$.
@@
IV.2 Numerical experiments
IV.3 Stability analysis for continuous encodings
@@
Encoder after the first transition
@@
$p(h|x)=p_{0}(h)\left[1+\frac{h}{\sigma^{2}}\phi_{1}(x)m_{1}+\frac{(h^{2}-\sigma^{2})}{2\sigma^{2}}\phi_{1}^{2}(x)m_{1}^{2}+\frac{h(h^{2}-3\sigma^{2})}{6\sigma^{2}}\phi_{1}^{3}(x)m_{1}^{3}+\frac{(h^{4}-6h^{2}\sigma^{2}+3\sigma^{4})}{24\sigma^{8}}\phi_{1}^{4}(x)m_{1}^{4}+\mathcal{O}(m_{1}^{5})\right]$
@@
$I(X,H)$ $=\int\text{d}x\,p(x)\frac{1}{2\sigma^{2}}\left(\phi_{1}(x)-\langle\phi_{1}\rangle\right)^{2}m_{1}^{2}+\frac{1}{4\sigma^{4}}\left(2(\phi_{1}(x)-\langle\phi_{1}\rangle)^{2}+\langle\phi_{1}\rangle^{2}-\langle\phi_{1}^{2}\rangle\right)(\langle\phi_{1}\rangle^{2}-\langle\phi_{1}^{2}\rangle)m_{1}^{4}$
@@
$=\frac{1}{4\sigma^{4}}\left(\left\langle\left(\langle\phi\rangle_{p(\cdot|y)}^{2}-\langle\phi^{2}\rangle_{p(\cdot|y)}\right)^{2}\right\rangle_{y}-\left(\langle\phi\rangle^{2}-\langle\phi^{2}\right\rangle)^{2}\right)m_{1}^{4}$
$=\left(\left\langle\left(\frac{\langle\phi\rangle_{p(\cdot|y)}^{2}-\langle\phi^{2}\rangle_{p(\cdot|y)}}{2\sigma^{2}}\right)^{2}\right\rangle_{y}-\left(\frac{\langle\phi\rangle^{2}-\langle\phi^{2}\rangle)}{2\sigma^{2}}\right)^{2}\right)m_{1}^{4}$
@@
$\mathcal{L}_{\text{IB}}$ $\propto 2\sigma^{2}(\text{Var}(\phi_{1})-\beta\text{Var}(\phi_{1}|y))m_{1}^{2}-\left((1-\beta)\text{Var}^{2}(\phi_{1})-\beta\text{Var}^{2}(\phi_{1}|y))\right)m_{1}^{4}$ (S76)
@@
IV.4 Limitations of the above results
@@
IV.5 “Distillation” of the optimal encoder
@@ V Variational IB
Equation (S96) is the normal form (in polar coordinates) for dynamics near a Hopf bifurcation [118]. For $\mu>0$ this system exhibits a circular limit cycle of radius $\sqrt{\mu}$ (Fig. S8b). In Section B we show that this system has infinite purely imaginary eigenvalues $\lambda_{n}=in\omega$ and that as a result, encoders which exactly encode the angle coordinate, $p(h|r,\theta)\propto\delta(h-\theta)$, are solutions of the IB optimization problem.
The exact IB calculation breaks down for perfectly deterministic dynamics, hence we add a small amount of white noise to the dynamics. While this slightly perturbs the spectral content of the transfer operator, we still find our results to be consistent with the those expected for the deterministic case (see Section B). As we increase the size of the encoding alphabet $N_{H}$, the encoder partitions space by finer and finer angular wedges. The learned encoding is invariant under permutations of the encoded symbols. Upon reordering, we find that, for large alphabets $N_{H}\gg 1$, the encoder indeed approximates $p(h|r,\theta)\propto\delta(h-\theta)$ as expected (Fig. S8c).
Using VIB, we learn a continuous $h_{t}$ which can be computed directly from samples of the state variable. The encoding $h_{t}$ learned by VIB closely approximates the angle coordinate (Fig. S8e-f) and is nearly uncorrelated with $r$ (inset). This shows that our mathematical results illustrated for exact IB in Sec. IV hold also in the approximate framework of VIB.
@@
IB is an inherently probabilistic framework. To handle deterministic dynamics, we inject noise by introducing a stochastic sampling scheme. This step is necessary only for deterministically simulated dynamics; in our experience experimental data is sufficiently noisy on its own. Concretely, rather than taking as our IB relevance variable $Y=X_{t+\Delta t}$, we take $Y=X_{t+\Delta t+\eta}$ where $\eta$ is a random uniformly-distributed time shift. Despite using a different relevance variable, this yields essentially the same optimal encoding as (S47), where the eigenvalue $\text{e}^{\lambda_{n}\Delta t}$ is replaced by $\int d\eta\,p(\eta)\exp(\lambda_{n}(\Delta t+\eta))$. Crucially, the encoder retains its dependence on the transfer operator eigenfunctions $\phi_{n}(x)$ as before. In general, this need not be the case: selecting a new relevance variable changes the IB objective and will generically lead to a different encoder. Our choice of stochasticity which we introduce to the dynamics is chosen in such a way to preserve the form of the encoder.
@@
VI VIB for fluid flow around a cylinder
@@
Here we provide some additional details on the computation used to generate components of Fig. 5 of the main text. The true Koopman modes were computed using dynamical mode decomposition (DMD) [15, 16, 120], also described below in Section X. DMD attempts to find a finite dimensional approximation of the Koopman operator using $n$ snapshots of the system’s state $x\in\mathbb{R}^{d}$ which are assembled into a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$. One then attempts to find a linear evolution operator $K$ which propagates the state forward in time
@@
where $\mathbf{m}^{(n)}$ is the $n$-th “Koopman mode” and angled brackets denote integration over space. To compute these modes and the corresponding transfer operator spectrum, we use dynamic mode decomposition (DMD; see SI Section VI) [15, 16].

### v4

@@ E.4 Summary
Figure S3: Information gain due to an eigenvalue pile-up. (a) To study the role of a gap closing compared to the pile-up of eigenvalues (beyond the dominant one) in the mutual information $I(X_{t};X_{t+\Delta t})$ we build a matrix $P$ with which we can tune the eigenvalues via a parameter $\varepsilon$. (b) Eigenvalues for the matrix shown in (a) with varying parameter values $\varepsilon$; for $\varepsilon\rightarrow 0.5$, the spectral gap $\lambda_{0}-\lambda_{1}$ closes. (c) Eigenvalues for the matrix shown in (a), but with all $\frac{1}{2}(\frac{1}{2}-\varepsilon)^{2}$ terms subtracted with a constant (here $1/8$). Here, the spectral gap $\lambda_{0}-\lambda_{1}$ does not close but there is an accumulation of eigenvalues at $\varepsilon\rightarrow 0.5$. (d-e) Mutual information for $\Delta t=1$ for the matrices in (b-c), respectively, as a function of $\varepsilon$. The curve from the other scenario is shown in gray for comparison. Both exhibit a clear peak in information which differs only slightly in magnitude, showing that the information peak in Fig. 2 of the main text is not due to the closing gap alone, but rather due to contributions from subdominant eigenvalues.
@@
$D_{\text{KL}}\left(p(x_{n+1}|x_{n},x_{n-1})\lVert p(x_{n+1}|x_{n})\right).$
@@
A system which is non-Markovian can often be made Markovian by introducing time-lagged variables.One constructs a new dynamical variable which is a concatenation of the previous states of the system,
@@
In Ref. [41] (and reproduced above in Eq. (46)), it is shown that the encoding minimizing the IB objective satisfies the implicit equation
@@
In the $D_{\text{KL}}$ in Eq. (S40), the integration is done with respect to the common random variable $X_{t+\Delta t}$ of the two conditional distributions.
@@
Note the appearance of the encoder $p(h_{t}|x_{t})$, making Eq. (S40) only an implicit solution of the optimization problem.
@@
The Kullback-Leibler divergence in Eq. (S40) can be represented as a sum of two terms,
@@
The first term has no dependence on $h_{t}$ and can hence be absorbed into the normalization $N(x)$. Plugging the decomposition Eq. (S44) into the second term leads directly to
@@
S4.1 Perturbation theory at the uniform encoder
@@
Theorem – Consider a transfer operator $U$ with infinitesimal generator $\mathcal{L}_{U}$. Assume that $\mathcal{L}_{U}$ has a discrete spectrum with eigenvalues satisfying $0=\text{Re }\lambda_{0}>\text{Re }\lambda_{1}>\text{Re }\lambda_{2}\gg\text{Re }\lambda_{3}$.... Then, for $\beta$ just above the first IB transition $\beta_{1}$ so that $f_{n}(h)\rightarrow 0$, the optimal encoder is given approximately by
@@
Proof - To show this we compute the $f_{n}(h)$ coefficients at the onset of instability. The instability at the first IB transition at $\beta=\beta_{1}$ corresponds to the emergence of negative eigenvalues in the Hessian of the IB loss,
@@
Angled brackets with no subscript correspond to an average with respect to the steady state distribution, $\langle\cdot\rangle=\int\text{d}x\,\cdot\,p(x)$. Written out more explicitly, the second term is
@@
which are increasing in order (remember $\lambda_{i}\leq 0$). It follows that at the first transition, only $f_{1}$ becomes non-zero, and hence the encoder takes the form given by Eq. (S39). In the equilibrium case, the encoder learns exclusively the first eigenfunction, with no correction due to the second eigenfunction.
@@
Similar to the equilibrium case, we see that at the first transition the encoder depends only on $f_{1}$, giving Eq. (S39). In contrast to the equilibrium case, there may be a small correction due to the second eigenfunction, however this becomes exponentially small for long times $\Delta t$.
@@
S4.2 Optimal IB encoder after the transition
We characterize the transition by expanding to fourth order and showing that it is continuous. We take
$p(h_{\mu}|x)=\frac{f_{0}^{\mu}\exp\left[\tilde{\phi}_{1}(x)f_{1}^{\mu}\right]}{\sum_{h_{\lambda}}f_{0}^{\lambda}\exp\left[\tilde{\phi}_{1}(x)f_{1}^{\lambda}\right]}$ (S61)
and expand to fourth order in $f_{1}^{\mu}$. We have
$p(h_{\mu}|x)$ $\approx f_{0}^{\mu}\Big[1+\tilde{\phi}_{1}(x)f_{1}^{\mu}$ (S62)
$\qquad+\frac{1}{2}\tilde{\phi}_{1}(x)^{2}\left((f_{1}^{\mu})^{2}-\langle f_{1}^{2}\rangle_{h}\right)$ (S63)
$\qquad+\frac{1}{6}\tilde{\phi}_{1}(x)^{3}\left((f_{1}^{\mu})^{3}-3f_{1}^{\mu}\langle f_{1}^{2}\rangle_{h}-\langle f_{1}^{3}\rangle_{h}\right)$ (S64)
$\qquad+\frac{1}{24}\tilde{\phi}_{1}(x)^{4}\left((f_{1}^{\mu})^{4}-4f_{1}^{\mu}\langle f_{1}^{3}\rangle_{h}-6(f_{1}^{\mu})^{2}\langle f_{1}^{2}\rangle_{h}+6\langle f_{1}^{2}\rangle_{h}^{2}-\langle f_{1}^{4}\rangle_{h}\right)\Big]$ (S65)
$\equiv f_{0}^{\mu}\Big[1+\tilde{\phi}_{1}(x)f_{1}^{\mu}+\frac{1}{2}\tilde{\phi}_{1}(x)^{2}A_{\mu}+\frac{1}{6}\tilde{\phi}_{1}(x)^{3}B_{\mu}+\frac{1}{24}\tilde{\phi}_{1}(x)^{4}C_{\mu}\Big]$ (S66)
where $\langle G\rangle_{h}=\sum_{\mu}f_{0}^{\mu}G(h_{\mu})$ denotes the average over all $h$ (recall that $p(h_{\mu})=f_{0}^{\mu}$). The expression for $p(h_{\mu})$ is similar, except all $\phi(x)$ are replaced with averages over $p(x)$: $\langle\phi\rangle_{x}$. Similarly, we have
$\log p(h_{\mu}|x)$ $=\log f_{0}^{\mu}+\tilde{\phi}_{1}(x)f_{1}^{\mu}-\frac{1}{2}\tilde{\phi}_{1}(x)^{2}\langle f_{1}^{2}\rangle_{h}-\frac{1}{6}\tilde{\phi}_{1}(x)^{3}\langle f_{1}^{3}\rangle_{h}-\frac{1}{24}\tilde{\phi}_{1}(x)^{4}\left(\langle f_{1}^{4}\rangle_{h}-3\langle f_{1}^{2}\rangle_{h}^{2}\right)$ (S67)
We can now readily compute the mutual information
$I(X_{t},H_{t})$ $=\sum_{\mu}\int\text{d}x\,p(h_{\mu}|x)p(x)\left[\log p(h_{\mu}|x)-\log p(h_{\mu})\right]$
$=\left[\langle\tilde{\phi}_{1}^{2}\rangle_{x}\langle f_{1}^{2}\rangle_{h}+\frac{1}{2}\langle\tilde{\phi}_{1}^{3}\rangle_{x}\langle f_{1}^{3}\rangle_{h}+\frac{1}{6}\langle\tilde{\phi}_{1}^{4}\rangle_{x}\left(\langle f_{1}^{4}\rangle_{h}-3\langle f_{1}^{2}\rangle_{h}^{2}\right)\right]$
$\qquad-\left[\frac{1}{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x}\langle f_{1}^{2}\rangle_{h}+\frac{1}{6}\langle\tilde{\phi}_{1}^{3}\rangle_{x}\langle f_{1}^{3}\rangle_{h}+\frac{1}{24}\langle\tilde{\phi}_{1}^{4}\rangle_{x}\left(\langle f_{1}^{4}\rangle_{h}-3\langle f_{1}^{2}\rangle_{h}^{2}\right)\right]$
$\qquad-\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}\left(\langle f_{1}^{4}\rangle_{h}-\langle f_{1}^{2}\rangle_{h}^{2}\right)$
$=\frac{1}{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x}\langle f_{1}^{2}\rangle_{h}+\frac{1}{3}\langle\tilde{\phi}_{1}^{3}\rangle_{x}\langle f_{1}^{3}\rangle_{h}+\frac{1}{8}\langle\tilde{\phi}_{1}^{4}\rangle_{x}\left(\langle f_{1}^{4}\rangle_{h}-3\langle f_{1}^{2}\rangle_{h}^{2}\right)-\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}\left(\langle f_{1}^{4}\rangle_{h}-\langle f_{1}^{2}\rangle_{h}^{2}\right).$ (S68)
The computation for the prediction term $I(H_{t},X_{t+\Delta t})$ proceeds similarly, if slightly messier:
$I(H_{t},Y)$ $=\frac{1}{2}\langle f_{1}^{2}\rangle_{h}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\right\rangle_{y}$
$\quad+\langle f_{1}^{3}\rangle_{h}\left[\frac{1}{2}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}-\frac{1}{6}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{3}\right\rangle_{y}\right]$
$\quad+\langle f_{1}^{4}\rangle_{h}\left[\frac{1}{6}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{3}\rangle_{x|y}\right\rangle_{y}+\frac{1}{8}\left\langle\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}^{2}\right\rangle_{y}-\frac{1}{4}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}-\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}+\frac{1}{12}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{4}\right\rangle_{y}\right]$
$\quad+\langle f_{1}^{2}\rangle_{h}^{2}\left[-\frac{1}{2}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{3}\rangle_{x|y}\right\rangle_{y}-\frac{1}{8}\left\langle\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}^{2}\right\rangle_{y}+\frac{1}{4}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}+\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}\right]$ (S69)
Together, these yield the full Lagrangian:
$\mathcal{L}_{\text{IB}}$ $=\frac{1}{2}\langle f_{1}^{2}\rangle_{h}\Big[\langle\tilde{\phi}_{1}^{2}\rangle_{x}-\beta\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\right\rangle_{y}\Big]$
$\quad+\langle f_{1}^{3}\rangle_{h}\left[\frac{1}{3}\langle\tilde{\phi}_{1}^{3}\rangle_{x}-\beta\left(\frac{1}{2}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}-\frac{1}{6}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{3}\right\rangle_{y}\right)\right]$
$\quad+\langle f_{1}^{4}\rangle_{h}\Bigg[\frac{1}{8}\left(\langle\tilde{\phi}_{1}^{4}\rangle_{x}-\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}\right)$
$\qquad\qquad\quad-\beta\left(\frac{1}{6}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{3}\rangle_{x|y}\right\rangle_{y}+\frac{1}{8}\left\langle\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}^{2}\right\rangle_{y}-\frac{1}{4}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}-\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}+\frac{1}{12}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{4}\right\rangle_{y}\right)\Bigg]$
$\quad+\langle f_{1}^{2}\rangle_{h}^{2}\Bigg[\frac{1}{8}\left(\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}-3\langle\tilde{\phi}_{1}^{4}\rangle_{x}\right)$
$\qquad\qquad\quad-\beta\left(-\frac{1}{2}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}\langle\tilde{\phi}_{1}^{3}\rangle_{x|y}\right\rangle_{y}-\frac{1}{8}\left\langle\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}^{2}\right\rangle_{y}+\frac{1}{4}\left\langle\langle\tilde{\phi}_{1}\rangle_{x|y}^{2}\langle\tilde{\phi}_{1}^{2}\rangle_{x|y}\right\rangle_{y}+\frac{1}{8}\langle\tilde{\phi}_{1}^{2}\rangle_{x}^{2}\right)\Bigg]$ (S70)
For further analysis, we consider the case where $N_{H}=2$. This is sufficient to study the behavior at the transition because more than one encoding degree of freedom would be redundant. We further consider the case of a symmetric perturbation, in which $f_{1}^{0}=-f_{1}^{1}\equiv f$. This simplifies the moments of $f$ significantly, yielding
$\langle f_{1}^{2}\rangle_{h}$ $=f^{2}$ (S71)
$\langle f_{1}^{3}\rangle_{h}$ $=0$ (S72)
$\langle f_{1}^{4}\rangle_{h}$ $=f^{4}$ (S73)
Then the IB loss simplifies dramatically to
$\mathcal{L}_{\text{IB}}$ $=\frac{1}{2}f^{2}\left[\langle\phi_{1}^{2}\rangle_{x}-\beta\left\langle\langle\phi_{1}\rangle_{x|y}^{2}\right\rangle_{y}\right]-\frac{1}{12}f^{4}\left[3\langle\phi_{1}^{4}\rangle_{x}-\beta\left(4\left\langle\langle\phi_{1}\rangle_{x|y}\langle\phi_{1}^{3}\rangle_{x|y}\right\rangle_{y}-\left\langle\langle\phi_{1}\rangle_{x|y}^{4}\right\rangle_{y}\right)\right]$ (S74)
which has minima at
$f^{*}=\pm\sqrt{\frac{3\left(\langle\phi_{1}^{2}\rangle_{x}-\beta\left\langle\langle\phi_{1}\rangle_{x|y}^{2}\right\rangle_{y}\right)}{3\langle\phi_{1}^{4}\rangle_{x}-\beta\left(4\left\langle\langle\phi_{1}\rangle_{x|y}\langle\phi_{1}^{3}\rangle_{x|y}\right\rangle_{y}-\left\langle\langle\phi_{1}\rangle_{x|y}^{4}\right\rangle_{y}\right)}}.$ (S75)
A stable solution emerges when
$\beta>\frac{\langle\phi_{1}^{2}\rangle_{x}}{\left\langle\langle\phi_{1}\rangle_{x|y}^{2}\right\rangle_{y}},$ (S76)
which matches the transition criterium for a Gaussian encoder (see below). Together, this shows that the value of $f$ changes continuously at the transition.
S4.3 Stability analysis for continuous encodings
@@
S4.4 Gaussian encoder after the first transition
@@
$p(h|x)$ $=p_{0}(h)\Big[1+\frac{h}{\sigma^{2}}\phi_{1}(x)m_{1}+\frac{(h^{2}-\sigma^{2})}{2\sigma^{2}}\phi_{1}^{2}(x)m_{1}^{2}+\frac{h(h^{2}-3\sigma^{2})}{6\sigma^{2}}\phi_{1}^{3}(x)m_{1}^{3}+$
$\quad\quad+\frac{(h^{4}-6h^{2}\sigma^{2}+3\sigma^{4})}{24\sigma^{8}}\phi_{1}^{4}(x)m_{1}^{4}+\mathcal{O}(m_{1}^{5})\Big]$
@@
$I(X,H)$ $=\int\text{d}x\,p(x)\frac{1}{2\sigma^{2}}\left(\phi_{1}(x)-\langle\phi_{1}\rangle\right)^{2}m_{1}^{2}+$
$\quad+\int\text{d}x\,p(x)\frac{1}{4\sigma^{4}}\left(2(\phi_{1}(x)-\langle\phi_{1}\rangle)^{2}+\langle\phi_{1}\rangle^{2}-\langle\phi_{1}^{2}\rangle\right)(\langle\phi_{1}\rangle^{2}-\langle\phi_{1}^{2}\rangle)m_{1}^{4}$
@@
$=\frac{1}{4\sigma^{4}}\left(\left\langle\left(\langle\phi\rangle_{p(\cdot|y)}^{2}-\langle\phi^{2}\rangle_{p(\cdot|y)}\right)^{2}\right\rangle_{y}-\left(\langle\phi\rangle^{2}-\langle\phi^{2}\rangle\right)^{2}\right)m_{1}^{4}$
$=\left(\left\langle\left(\frac{\langle\phi\rangle_{p(\cdot|y)}^{2}-\langle\phi^{2}\rangle_{p(\cdot|y)}}{2\sigma^{2}}\right)^{2}\right\rangle_{y}-\left(\frac{\langle\phi\rangle^{2}-\langle\phi^{2}\rangle}{2\sigma^{2}}\right)^{2}\right)m_{1}^{4}$
@@
$\mathcal{L}_{\text{IB}}$ $\propto 2\sigma^{2}(\text{Var}(\phi_{1})-\beta\langle\text{Var}(\phi_{1}|y)\rangle_{y})m_{1}^{2}-\left((1-\beta)\text{Var}^{2}(\phi_{1})-\beta\langle\text{Var}^{2}(\phi_{1}|y)\rangle_{y}\right)m_{1}^{4}$ (S84)
@@
S4.5 Stability analysis of $I(H_{t},H_{t+\Delta t})$
$=p(h_{\nu})(\delta_{\mu\nu}-p(h_{\mu}))\delta_{n0}$ (S87)
$=\int\text{d}x_{0}\text{d}x_{t}\,p(h_{t}|x_{t})p(h_{0}|x_{0})p(x_{0},x_{t})$
Evaluated at the uniform encoder, the first term disappears because we have the derivative of a constant. From the equations for the derivatives Eq. (S86) and Eq. (S87), we can conclude $\partial_{n}^{\nu}\partial_{m}^{\mu}I(H_{0},H_{t})\propto\delta_{n0}\delta_{m0}$, which shows that we lose all information about the eigenfunctions. The Hessian of the Lagrangian for IB modified by the term $I(H_{0},H_{t})$ will therefore have the same stability properties as the original IB problem.
S4.6 Numerical experiments
@@
S4.7 Limitations of the above results
@@
S4.8 “Distillation” of the optimal encoder
@@
Equation (S107) is the normal form (in polar coordinates) for dynamics near a Hopf bifurcation [155]. For $\mu>0$ this system exhibits a circular limit cycle of radius $\sqrt{\mu}$ (Fig. S8b). In the following we show that this system has infinite purely imaginary eigenvalues $\lambda_{n}=in\omega$ and that as a result, encoders which exactly encode the angle coordinate, $p(h|r,\theta)\propto\delta(h-\theta)$, are solutions of the IB optimization problem.
The exact IB calculation breaks down for perfectly deterministic dynamics, hence we add a small amount of white noise to the dynamics. While this slightly perturbs the spectral content of the transfer operator, we still find our results to be consistent with the those expected for the deterministic case (see the next section). As we increase the size of the encoding alphabet $N_{H}$, the encoder partitions space by finer and finer angular wedges. The learned encoding is invariant under permutations of the encoded symbols. Upon reordering, we find that, for large alphabets $N_{H}\gg 1$, the encoder indeed approximates $p(h|r,\theta)\propto\delta(h-\theta)$ as expected (Fig. S8c).
Using VIB, we learn a continuous $h_{t}$ which can be computed directly from samples of the state variable. The encoding $h_{t}$ learned by VIB closely approximates the angle coordinate (Fig. S8e-f) and is nearly uncorrelated with $r$ (inset). This shows that our mathematical results illustrated for exact IB in Sec. S4 hold also in the approximate framework of VIB.
Optimal encoding for Hopf normal form dynamics
$\dot{x}$ $=(\mu-x^{2}-y^{2})x-\omega y$ (S108)
$\dot{y}$ $=(\mu-x^{2}-y^{2})y-\omega x$ (S109)
Expressed in polar coordinates, we have
$=r(\mu-r^{2})(\partial_{r}\phi)+\omega\partial_{\theta}\phi$ (S113)
For the angle coordinate alone the situation is much simpler,
$\omega\partial_{\theta}\phi(\theta)=\lambda\phi(\theta)$
$\phi_{n}(\theta)=e^{in\theta},\quad\quad\lambda_{n}=in\omega.$
We next ask whether an encoder of the form $p(h|\theta)\propto\delta(h-\theta)$ is a solution. Recall that the probability distribution appearing in Eq. (S116) is a distribution over future positions, $p(R_{t+\Delta t}=r^{\prime},\Theta_{t+\Delta t}=\theta+\omega\Delta t|h)$. This distribution can be calculated as
$\propto\delta((\theta_{t+\Delta t}-\omega\Delta t)-h)$
where we used that the two terms in the integrand of the first equation are both delta functions. Plugging this into Eq. (S116) shows that $p(h|\theta)\propto\delta(h-\theta)$ is consistent, and hence is an optimal encoding. In the presence of noise, the eigenvectors are perturbed and may gain a dependence on $r$ (see Fig. S9).
@@
IB is an inherently probabilistic framework. To handle deterministic dynamics, we inject noise by introducing a stochastic sampling scheme. This step is necessary only for deterministically simulated dynamics; in our experience experimental data is sufficiently noisy on its own. Concretely, rather than taking as our IB relevance variable $Y=X_{t+\Delta t}$, we take $Y=X_{t+\Delta t+\eta}$ where $\eta$ is a random uniformly-distributed time shift. Despite using a different relevance variable, this yields essentially the same optimal encoding as Eq. (S39), where the eigenvalue $\text{e}^{\lambda_{n}\Delta t}$ is replaced by $\int d\eta\,p(\eta)\exp(\lambda_{n}(\Delta t+\eta))$. Crucially, the encoder retains its dependence on the transfer operator eigenfunctions $\phi_{n}(x)$ as before. In general, this need not be the case: selecting a new relevance variable changes the IB objective and will generically lead to a different encoder. Our choice of stochasticity which we introduce to the dynamics is chosen in such a way to preserve the form of the encoder.
@@
S6 VIB for fluid flow around a cylinder
@@
Here we provide some additional details on the computation used to generate components of Fig. 5 of the main text. The true Koopman modes were computed using dynamical mode decomposition (DMD) [21, 22, 157], also described below in Section S10. DMD attempts to find a finite dimensional approximation of the Koopman operator using $n$ snapshots of the system’s state $x\in\mathbb{R}^{d}$ which are assembled into a data matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$. One then attempts to find a linear evolution operator $K$ which propagates the state forward in time
@@
where $\mathbf{m}^{(n)}$ is the $n$-th “Koopman mode” and angled brackets denote integration over space. To compute these modes and the corresponding transfer operator spectrum, we use dynamic mode decomposition (DMD; see SI Section S6) [21, 22].